An experiment consists of rolling one fair number cube. Find the probability of each event
P(3)
P(7)
P(<5)
P(2 or odd)
Rolling two fair number cubes
P(total shown =9)
P(total shown < or = 12)

Answer by edjones(8007)

P(3)=1/6
P(7)=0
P(<5)=4/6=2/3
P(2 or odd)=4/6=2/3
Rolling two fair number cubes
P(total shown =9)=45,54,63,36=4/36=1/9
P(total shown < or = 12)35/36 (6,6 is 12 All other combinations are lower)
.
Ed